OpenStudy (anonymous):

@ HASHIR EXPAND THIS FOR ME (x+3)^-2

OpenStudy (anonymous):

1/x^2+6x+9

OpenStudy (anonymous):

how did u get that explain

OpenStudy (anonymous):

idk u explain i solved it using my big bros help

OpenStudy (anonymous):

x^(3/2) = 3^3 x = 3^[3(2/3)] x = 3^2 x= 9

OpenStudy (anonymous):

i m in kindergarten

OpenStudy (anonymous):

(x+3)^-2 = (x^2 + 6x + 9)^-1 1/x^2+6x+9

OpenStudy (anonymous):

1/(x^2+6x+9)

OpenStudy (anonymous):

too hard for my comprehension hashir

OpenStudy (anonymous):

(x+3)^-2=1/((x+3)^2)

OpenStudy (anonymous):

ah I wanted to do this this is not fair : ( haha joking :D